### v4

@@ Abstract
This paper investigates preconditioning techniques for solving kernel ridge regression (KRR) problems with a medium to large number of data points ($10^{4}\leq N\leq 10^{7}$), and it introduces two new methods with state-of-the-art performance. The first method, RPCholesky preconditioning, accurately solves the full-data KRR problem in $\mathcal{O}(N^{2})$ arithmetic operations, assuming sufficiently rapid polynomial decay of the kernel matrix eigenvalues. The second method, KRILL preconditioning, offers an accurate solution to a restricted version of the KRR problem involving $k\ll N$ selected data centers at a cost of $\mathcal{O}((N+k^{2})k\log k)$ operations. The proposed methods efficiently solve a broad range of KRR problems, making them ideal for practical applications.
@@ 1 Motivation
Kernel ridge regression (KRR) is a machine learning method for learning an unknown function $y=f(\bm{x})$ from a set of input–output pairs
@@
that models the input–output relationship. KRR is widely used in data science and scientific computing, for example, in predicting the chemical properties of molecules [18, 54]. Section 2 provides more details about KRR.
@@
In practice, KRR is often limited to small- or medium-sized data sets ($N\leq 10^{4}$) because the computation time can grow rapidly with the number of data points. KRR involves linear algebra computations with the $N\times N$ kernel matrix $\bm{A}$ with entries $a_{ij}=K(\bm{x}^{(i)},\bm{x}^{(j)})$. To obtain the prediction function $\widehat{f}$, we must solve an $N\times N$ positive-definite linear system to high accuracy, which requires $\mathcal{O}(N^{3})$ arithmetic operations using the standard direct method based on Cholesky decomposition. Because of this poor computational scaling, survey articles on machine learning for scientific audiences (such as [51]) suggest using KRR for small data sets and neural networks for larger data sets. Reliably scaling KRR to larger data sets with $N\geq 10^{4}$ remains an open research challenge.
To improve the scalability of KRR, two main approaches have been developed. The first approach aims to solve the full-data KRR problem accurately in $\mathcal{O}(N^{2})$ operations using preconditioned conjugate gradient (CG) [16, 5]. The second approach restricts the KRR problem to a small number of $k\ll N$ centers chosen from the data points [48] and aims to solve the restricted problem in $\mathcal{O}(kN)$ operations using preconditioned CG [44].
@@
However, due to the ill-conditioned nature of many KRR problems, CG-based approaches can only succeed when there is a high-quality preconditioner. Consequently, much of the research on solving KRR problems over the past decade has focused on proposing preconditioners [16, 44, 23, 45, 22, 47], empirically testing preconditioners [16, 47, 8], and theoretically analyzing preconditioners [5, 44, 23, 45, 22]. Despite this extensive body of work, a user of KRR might struggle to identify the best preconditioner, since many ideas have been proposed and there has not been a recent comparison of the available approaches.
@@
The primary purpose of a preconditioner is to minimize the number of CG iterations required to solve a linear system to a desired level of accuracy. For KRR, each CG iteration can be computationally expensive as it involves either one multiplication with the $N\times N$ kernel matrix (for full-data KRR) or two multiplications with the $N\times k$ kernel submatrix (for restricted KRR). Additionally, the kernel matrix or submatrix may not fit in working memory, in which case kernel matrix entries must be accessed from storage or regenerated from data at each CG iteration, which adds to the computational cost. This paper seeks to reduce the expense of KRR by identifying preconditioners that are both efficient to construct and capable of controlling the number of CG iterations required for convergence.
@@
To be useful in practice, a preconditioner must work reliably for a range of KRR instances, and it must be robust against adverse conditions (such as poorly conditioned kernel matrices). Additionally, the preconditioner should not require any delicate parameter tuning. Such a robust and no-hassle preconditioner is vital for general-purpose scientific software that can handle a variety of KRR problems framed by different users. Additionally, such a preconditioner is critical for cross-validation, as it enables searching for the best kernel model while having confidence that the preconditioner can handle the resulting linear systems efficiently. Unfortunately, as this paper shows, many existing KRR preconditioners are prone to failure when applied to specific KRR problems, limiting their range of applicability.
@@
To meet the need for higher-quality preconditioners, this paper harnesses randomized matrix approximation schemes [12, 15] that have not been previously applied in the context of KRR but are valuable for achieving robustness. We use these approximations to devise two new randomized preconditioners that allow us to solve more KRR problems in fewer iterations than the techniques in current use [16, 44, 23, 45, 22]. The RPCholesky preconditioner (Algorithm 1) can be applied to the full-data KRR problem, while the KRILL preconditioner (Algorithm 2) can be applied to the restricted KRR problem involving $k\ll N$ data centers. Both preconditioners are cheap to implement, are effective for a wide range of KRR problems, and are supported by strong theoretical guarantees.
@@
The RPCholesky preconditioner is designed to solve linear systems involving the full-data $N\times N$ kernel matrix. The preconditioning algorithm constructs a nearly optimal low-rank approximation of the kernel matrix [12], and the preconditioner employs a shift of this low-rank approximation. Under favorable eigenvalue decay conditions, RPCholesky preconditioning brings the number of CG iterations down to acceptable levels ($10^{0}$$10^{2}$). Yet it is important to recognize that any preconditioner based on low-rank approximation, including RPCholesky, requires that the rank of the kernel approximation is at least as high as the number of large eigenvalues in the kernel matrix. Thankfully, many problems in scientific machine learning do exhibit sufficient eigenvalue decay for this approach to be successful: see Section 4.1 for a case study drawn from quantum chemistry.
@@
The KRILL preconditioner is designed to solve restricted KRR problems (Section 2.2), which are linear systems involving an $N\times k$ submatrix of the full $N\times N$ kernel matrix. KRILL constructs a preconditioner by performing dimensionality reduction on the $N\times k$ kernel submatrix using a sparse random sign embedding [15]. KRILL is consistently successful in numerical tests, as it reduces the number of CG iterations to acceptable levels ($10^{0}$$10^{2}$) regardless of the kernel matrix eigenvalue decay. Due to this reliability and robustness, just as krill are the foundation for the marine ecosystem, KRILL can serve as the foundation for large-scale KRR applications in the future
@@ 1.1 Plan for paper
The rest of this paper is organized as follows. Section 2 presents the RPCholesky and KRILL preconditioning strategies; Section 3 compares RPCholesky and KRILL to other preconditioners; Section 4 applies the RPCholesky and KRILL preconditioners to benchmark problems in computational chemistry and physics; Section 5 contains proofs of the theoretical results; and Section 6 offers conclusions.
@@ 1.2 Notation
[…]
@@ 2 Algorithms and best practices
For the KRR user’s convenience, this early section highlights our recommended approaches for solving KRR problems. Specifically, Section 2.1 introduces the RPCholesky preconditioner for solving the full-data KRR problem, while Section 2.2 introduces the KRILL preconditioner for solving the restricted KRR problem.
@@ 2.1 Full-data kernel ridge regression
$\widehat{f}(\bm{x};\bm{\beta})=\sum\nolimits_{i=1}^{N}\beta_{i}K(\bm{x}^{(i)},\bm{x}).$
The coefficients $\bm{\beta}\in\mathbb{R}^{N}$ for the function are chosen to minimize the regularized least-squares loss function
@@
The linear system can be solved in $\mathcal{O}(N^{3})$ time using a Cholesky decomposition of $\bm{A}+\mu\mathbf{I}$ and two triangular solves. However, if there is a medium or large number of data points ($N\geq 10^{4}$), we instead recommend solving (2.1) at a reduced cost using conjugate gradient with RPCholesky preconditioning, described below. We will prove that RPCholesky solves the full-data KRR problem in $\mathcal{O}(N^{2})$ time, provided the kernel matrix eigenvalues decay at a sufficiently fast polynomial rate (Theorem 2.2).
@@ 2.1.1 RPCholesky preconditioning
1:Positive semidefinite matrix $\bm{A}\in{}^{N\times N}$, right-hand-side vector $\bm{y}\in{}^{N}$, regularization coefficient $\mu$, approximation rank $r$, and tolerance $\varepsilon$.
@@
3:$\bm{F}\leftarrow\textsc{RPCholesky}(\bm{A},r,\min(100,r/10))$. $\triangleright$ See Algorithm 4.
4:$(\bm{U},\bm{\Sigma},\sim)\leftarrow\textsc{EconomySizeSVD}(\bm{F})$. $\triangleright$ $\bm{\widehat{A}}=\bm{U}\bm{\Sigma}^{2}\bm{U}^{*}$.
@@
6:$\bm{\beta}_{\star}\leftarrow\textsc{PCG}(\texttt{Product},\bm{y},\varepsilon,\texttt{Preconditioner})$. $\triangleright$ See Algorithm 3.
@@
There are many ways to obtain a low-rank approximation for this purpose, and we adopt the recently popular RPCholesky algorithm (Algorithm 4, analyzed in [12]).
The RPCholesky algorithm is a variant of partial Cholesky decomposition that chooses random pivots at each step according to an evolving probability distribution. For an input parameter $r\in\mathbb{N}$, the algorithm returns a rank-$r$ approximation in factorized form:
$\bm{\widehat{A}}=\bm{F}\bm{F}^{\ast}\quad\text{where}\quad\bm{F}\in{}^{N\times r}.$
The matrix $\bm{\widehat{A}}$ is random because it depends on the choice of random pivots. On average, $\bm{\widehat{A}}$ is comparable to the best rank-$r$ approximation of $\bm{A}$, as established in [12, Thm. 3.1]. The algorithm accesses only $(r+1)N$ entries of the kernel matrix; it uses $\mathcal{O}(rN)$ storage; and it expends $\mathcal{O}(r^{2}N)$ arithmetic operations.
We propose to select the approximation rank $r=\mathcal{O}(\sqrt{N})$, which ensures that we can obtain an eigenvalue decomposition of the preconditioner $\bm{P}$ in $\mathcal{O}(N^{2})$ operations. We then apply preconditioned CG (Algorithm 3, described in [24, §10.3]) to solve the KRR problem. Provided that $r=\mathcal{O}(\sqrt{N})$ suffices to obtain a good preconditioner, CG terminates in a constant number of iterations and the total operation count for RPCholesky preconditioning is $\mathcal{O}(N^{2})$. The pseudocode for RPCholesky preconditioning is given in Algorithm 1.
@@ 2.1.2 Empirical performance
Next we test the empirical performance of RPCholesky and four other preconditioners on the 20 regression and classification problems described in Table 1 of Appendix A. For each problem, we first randomly subsample $N=1.5\times 10^{4}$ data points. We standardize the features (subtract the mean, divide by the standard deviation) and measure similarity using the squared exponential kernel
@@
$K(\bm{x},\bm{y})=\exp\bigl{(}-\tfrac{1}{2\sigma^{2}}\lVert\bm{x}-\bm{y}\rVert^{2}\bigr{)}\quad\text{with bandwidth $\sigma=3$.}$ (2.3)
@@
We then formulate each KRR problem (2.1) with a small regularization parameter $\mu/N=10^{-7}$ (left panel) or a very small regularization parameter $\mu/N=10^{-10}$ (right panel) and choose the rank $r$ of the preconditioner to be $r=1000$. This choice of parameters is typical for full-data KRR problems [5, 22], and the same parameters are applied uniformly over the 20 regression and classification problems without any additional tuning. Note the smaller value of $\mu$ makes many of these problems highly ill-conditioned. We run 250 CG iterations and declare each KRR instance to be “solved” as soon as the relative residual $\lVert(\bm{A}+\mu\mathbf{I})\bm{\beta}-\bm{y}\rVert/\lVert\bm{\beta}\rVert$ falls below an error tolerance of $\varepsilon=10^{-3}$. This choice of $\varepsilon$ is justified by the fact that the test error plateaus after reaching this tolerance for all of the problems we considered.
Small regularization: $\mu=10^{-7}N$
Very small regularization: $\mu=10^{-10}N$
@@
Figure 1: Fraction of solved problems versus number of CG iterations for the 20 KRR problem instances in Table 1.
@@
Figure 1 charts the fraction of solved problems when using RPCholesky and four other preconditioners based on low-rank approximation. Two alternative pivoting schemes for partial Cholesky decomposition are greedy pivot selection [23, 53] and uniformly random pivot selection [16, 22]. Two additional methods for randomized low-rank approximation are ridge leverage score sampling (RLS, [1, 36, 45]) and random Fourier features (RFF, [16, 5]). For details of these methods, see Section 3.1.
@@
RPCholesky is the top-performing method in Figure 1, both when the regularization is small (left) and when the regularization is very small (right). The closest competitor method is RLS, yet RPCholesky consistently performs as well or better than RLS. For example, when $\mu/N=10^{-7}$, RPCholesky and RLS lead to nearly the same number of CG iterations (±$\pm 20\%$ differences) on 14 of the 20 test problems. On the remaining 6 problems, RPCholesky converges 1.2×1.2$\times$1.2 ×2.4×2.4$\times$2.4 × more quickly than RLS. As another advantage, RPCholesky is simpler to implement and faster to run than RLS, as it eliminates the cost of approximating the ridge leverage scores (see Section 3.1).
@@
We might worry that RPCholesky changes behavior depending on the random seed, but empirically the algorithm produces consistent results with just ±$\pm 5\%$ random variation in the number of CG iterations. To illustrate this point, Figure 2 shows the distribution of the relative residual error $\lVert(\bm{A}+\mu\mathbf{I})\bm{\beta}-\bm{y}\rVert/\lVert\bm{\beta}\rVert$ for the two problems with the fastest (COMET_MC_SAMPLE) and the slowest (sensorless) CG convergence. The lines indicate the median error and the shaded regions indicate the $20\%$-$80\%$ error quantiles, which are computed over 100 random runs. The 20%–80% quantiles are hardly distinguishable from the median lines, and they are even tighter for RPCholesky than the other stochastic methods (uniform, RLS, and RFF).
@@
sensorless Refer to caption
Figure 2: Relative residual versus number of CG iterations for the problem with the fastest (COMET_MC_SAMPLE, left) and slowest (sensorless, right) CG convergence when $\mu/N=10^{-7}$. Note the different vertical axis scales.
The importance of eigenvalue decay is evident from Figure 2. In the problem with rapid eigenvalue decay (COMET_MC_SAMPLE, left), RPCholesky converges to the accuracy threshold $\varepsilon=10^{-3}$ in just a single iteration. In the problem with slow eigenvalue decay (sensorless, right), RPCholesky converges 200×200$\times$200 × more slowly. All the methods based on low-rank approximation struggle to control the condition number for the sensorless problem, given that the eigenvalues of the kernel matrix decay slowly over 6 orders of magnitude. The dependence on eigenvalue decay is the main limitation of any preconditioner based on low-rank approximation. Nonetheless, even for this difficult problem, RPCholesky and the other preconditioners significantly improve on unpreconditioned CG.

### v5

@@ Abstract
This paper investigates preconditioned conjugate gradient techniques for solving kernel ridge regression (KRR) problems with a medium to large number of data points ($10^{4}\leq N\leq 10^{7}$), and it describes two methods with the strongest guarantees available. The first method, RPCholesky preconditioning, accurately solves the full-data KRR problem in $\mathcal{O}(N^{2})$ arithmetic operations, assuming sufficiently rapid polynomial decay of the kernel matrix eigenvalues. The second method, KRILL preconditioning, offers an accurate solution to a restricted version of the KRR problem involving $k\ll N$ selected data centers at a cost of $\mathcal{O}((N+k^{2})k\log k)$ operations. The proposed methods efficiently solve a range of KRR problems, making them well-suited for practical applications.
@@ 1 Motivation
Kernel ridge regression (KRR) is a machine learning method for approximating an unknown function $y=f(\bm{x})$ from a set of input–output pairs
@@
that models the input–output relationship. KRR is widely used in data science and scientific computing, for example, when predicting the chemical properties of molecules [20, 58]. Section˜2 provides more details about KRR.
@@
In practice, KRR is often limited to small- or medium-sized data sets ($N\leq 10^{4}$) because the computation time can grow rapidly with the number of data points. KRR involves linear algebra computations using the $N\times N$ kernel matrix $\bm{A}$ with entries $a_{ij}=K(\bm{x}^{(i)},\bm{x}^{(j)})$. To obtain the prediction function $\widehat{f}$, we must solve an $N\times N$ positive-definite linear system, which requires $\mathcal{O}(N^{3})$ arithmetic operations using the standard direct method based on Cholesky decomposition. Because of this poor computational scaling, survey articles on machine learning for scientific audiences, such as [55], suggest using KRR for small data sets and neural networks for larger data sets. Reliably scaling KRR to larger data sets with $N\geq 10^{4}$ data points remains an open research challenge.
@@
To improve the scalability of KRR, two approaches based on preconditioned conjugate gradient (CG) have been developed. The first approach aims to solve the full-data KRR problem in $\mathcal{O}(N^{2})$ operations [17, 5]. The second approach restricts the KRR problem to $k\ll N$ centers chosen from the data points [52] and aims to solve the restricted problem in $\mathcal{O}(kN)$ operations [47]. However, due to the ill-conditioned nature of many KRR problems, CG-based approaches can only succeed when there is a high-quality preconditioner. Much of the research on solving KRR problems over the past decade has focused on proposing preconditioners [17, 47, 26, 48, 25, 51], empirically testing preconditioners [17, 51, 8], and theoretically analyzing preconditioners [5, 47, 26, 48, 25]. Despite this extensive body of work, a user of KRR might struggle to identify the best preconditioner, since many ideas have been proposed and there has not been a recent comparison.
@@
The primary purpose of a preconditioner is to minimize the number of CG iterations required to solve a linear system to a desired level of accuracy. Each CG iteration can be computationally expensive as it involves either one multiplication with the $N\times N$ kernel matrix (for full-data KRR) or two multiplications with the $N\times k$ kernel submatrix (for restricted KRR). Additionally, the kernel matrix or submatrix may not fit in working memory, in which case kernel matrix entries must be accessed from storage or regenerated on the fly at each CG iteration, which adds to the computational cost. This paper seeks to reduce the expense of KRR by identifying preconditioners that are both efficient to construct and capable of controlling the number of CG iterations required for convergence.
@@
To be useful in practice, a preconditioner must work reliably for a range of KRR instances, and it must be robust against adverse conditions (such as poorly conditioned kernel matrices). Additionally, the preconditioner must not require delicate parameter tuning. Such a robust and no-hassle preconditioner is vital for scientific software that can handle KRR problems framed by different users. Such a preconditioner is critical for cross-validation, as it enables searching for the best kernel model while having confidence that the preconditioner can handle the resulting linear systems efficiently. Unfortunately, as this paper shows, many existing KRR preconditioners are prone to failure when applied to specific KRR problems, limiting their applicability.
@@
To meet the need for higher-quality preconditioners, this paper harnesses randomized matrix approximations [13, 16] that have not been previously applied in the context of KRR but are valuable for achieving robustness. The approximations lead to two randomized preconditioners that allow us to solve more KRR problems in fewer iterations than the techniques in current use [17, 47, 26, 48, 25]. The RPCholesky preconditioner (Algorithm˜1) can be applied to the full-data KRR problem, while the KRILL preconditioner (Algorithm˜2) can be applied to the restricted KRR problem involving $k\ll N$ data centers. Both preconditioners are cheap to implement, are effective for a range of KRR problems, and are supported by strong theoretical guarantees.
RPCholesky.
@@
The RPCholesky preconditioner is designed to solve linear systems involving the full-data $N\times N$ kernel matrix. The RPCholesky algorithm generates a low-rank approximation of the kernel matrix [13], and the preconditioner employs a shift of this low-rank approximation. Any preconditioner based on low-rank approximation, such as RPCholesky, can only be successful when the number of large eigenvalues in the kernel matrix is within the approximation rank. Yet thankfully, many problems in scientific machine learning do exhibit sufficient eigenvalue decay: see Section˜4.1 for a case study drawn from quantum chemistry. Under favorable eigenvalue decay conditions, RPCholesky preconditioning brings the number of CG iterations down to acceptable levels ($10^{0}$ to $10^{2}$).
KRILL.
@@
The KRILL preconditioner is designed to solve restricted KRR problems (Section˜2.2), which are linear systems involving an $N\times k$ submatrix of the full $N\times N$ kernel matrix. KRILL constructs a preconditioner by performing dimensionality reduction on the $N\times k$ kernel submatrix with a sparse random sign embedding [16]. KRILL always reduces the number of CG iterations to acceptable levels ($10^{0}$ to $10^{2}$) regardless of the kernel matrix eigenvalue decay. Thus, KRILL can serve as a robust and reliable tool for large-scale KRR applications, just as krill are a robust and reliable food source in the marine ecosystem.
Broader context.
For decades, preconditioned CG has been the method of choice for solving large-scale linear systems [49] and KRR problems [17, 47, 26, 48, 25]. However, emerging research suggests that row access methods can outperform preconditioned CG for select problems [44, 31]. The relative merits of preconditioned CG versus row access methods are being debated [33, 19], and so far there has not been a complete and definitive comparison between these linear solvers, in part because it is unclear which preconditioned CG and row access methods are most effective. The present work represents partial progress toward a resolution. It demonstrates that RPCholesky and KRILL have the strongest guarantees of any KRR preconditioners; therefore, these methods can serve as exemplars of the preconditioned CG approach in future comparisons.
@@ 1.1 Plan for paper
The rest of this paper is organized as follows. Section˜2 presents the RPCholesky and KRILL preconditioning strategies; Section˜3 compares RPCholesky and KRILL to other preconditioners; Section˜4 applies RPCholesky and KRILL to benchmark problems in computational chemistry and physics; Section˜5 presents proofs of the theoretical results; and Section˜6 offers conclusions.
@@ 1.2 Notation
For simplicity, we focus on the real setting, although our work extends to complex-valued kernels without significant modification. The transpose and Moore–Penrose pseudoinverse of $\bm{A}$ are denoted $\bm{A}^{*}$ and $\bm{A}^{\dagger}$. Double bars $\left\|\cdot\right\|$ indicate the Euclidean norm of a vector or the spectral norm of a matrix. The matrix condition number is $\kappa(\bm{A})\coloneqq\lVert\bm{A}\rVert\lVert\bm{A}^{-1}\rVert$. For a positive-definite matrix $\bm{M}$, the $\bm{M}$-weighted inner product norm is denoted $\lVert\bm{z}\rVert_{\bm{M}}\coloneqq(\bm{z}^{*}\bm{M}\bm{z})^{1/2}$. The function $\lambda_{i}(\bm{M})$ outputs the $i$th largest eigenvalue of $\bm{M}.$ The $\ell_{1}$ norm of the vector $\bm{x}$ is denoted by $\|\bm{x}\|_{\ell_{1}}=\sum_{i}|x_{i}|$. The symbol $|\mathsf{S}|$ denotes the cardinality of a set $\mathsf{S}$.
@@ 2 Algorithms and best practices
For the KRR user’s convenience, this early section highlights our recommended approaches for solving KRR problems. Specifically, Section˜2.1 introduces the RPCholesky preconditioner for solving the full-data KRR problem, while Section˜2.2 introduces the KRILL preconditioner for solving the restricted KRR problem.
@@ 2.1 Full-data kernel ridge regression
$\widehat{f}(\bm{x};\bm{\beta})=\sum_{i=1}^{N}\beta_{i}K(\bm{x}^{(i)},\bm{x}).$
The coefficients $\bm{\beta}\in\mathbb{R}^{N}$ for the function are chosen to minimize the regularized least-squares loss
@@
The linear system can be solved in $\mathcal{O}(N^{3})$ time using a Cholesky decomposition of $\bm{A}+\mu\mathbf{I}$ and two triangular solves. However, when there is a medium or large number of data points ($N\geq 10^{4}$), we instead recommend solving (2.1) at a reduced cost using conjugate gradient with RPCholesky preconditioning, described below. We will prove that the strategy based on RPCholesky solves the full-data KRR problem in $\mathcal{O}(N^{2})$ time, provided the kernel matrix eigenvalues decay at a sufficiently fast polynomial rate (Theorem˜2.2).
@@ 2.1.1 RPCholesky preconditioning
1:Positive semidefinite matrix $\bm{A}\in\real^{N\times N}$, right-hand-side vector $\bm{y}\in\real^{N}$, regularization coefficient $\mu$, approximation rank $r$, and tolerance $\varepsilon$.
@@
[…]
@@
[…]
@@
There are many ways to obtain a low-rank approximation for this purpose, and we adopt the RPCholesky algorithm (Algorithm˜4, analyzed in [13, 24]).
RPCholesky is a variant of the partial Cholesky decomposition that chooses random pivots at each step according to an evolving probability distribution. For an input parameter $r\in\mathbb{N}$, the algorithm returns a rank-$r$ approximation in factored form:
$\bm{\widehat{A}}=\bm{F}\bm{F}^{\ast}\quad\text{where}\quad\bm{F}\in\real^{N\times r}.$
The matrix $\bm{\widehat{A}}$ is random because it depends on the choice of random pivots. The average error of $\bm{\widehat{A}}$ is comparable to the best rank-$r$ approximation error, as established in [13, Thm. 3.1] and [24, Thm. 4.2]. The algorithm accesses only $(r+1)N$ entries of the kernel matrix; it uses $\mathcal{O}(rN)$ storage; and it expends $\mathcal{O}(r^{2}N)$ arithmetic operations.
We propose an approximation rank $r=\mathcal{O}(\sqrt{N})$, which ensures that we can obtain an eigenvalue decomposition of the preconditioner in $\mathcal{O}(N^{2})$ operations. We then apply preconditioned CG (Algorithm˜3, described in [27, §10.3]) to solve the KRR problem. If the rank $r=\mathcal{O}(\sqrt{N})$ suffices to obtain a good preconditioner, CG terminates in a constant number of iterations and the total operation count for RPCholesky preconditioning is $\mathcal{O}(N^{2})$. The pseudocode for RPCholesky preconditioning is given in Algorithm˜1.
@@ 2.1.2 Empirical performance
Here we test the empirical performance of RPCholesky and four other preconditioners across 20 regression and classification problems described in Table˜1 of Appendix˜A. For each problem, we randomly subsample $N=1.5\times 10^{4}$ data points. We standardize the features (subtract the mean, divide by the standard deviation) and measure similarity using the squared exponential kernel
@@
$K(\bm{x},\bm{y})=\exp\bigl(-\tfrac{1}{2\sigma^{2}}\lVert\bm{x}-\bm{y}\rVert^{2}\bigr)\quad\text{with bandwidth $\sigma=\sqrt{d}$.}$ (2.3)
@@
We then formulate each KRR problem (2.1) with a larger regularization parameter $\mu/N=10^{-7}$ (left panel) or a smaller regularization parameter $\mu/N=10^{-8}$ (right panel) and choose the rank $r$ of the preconditioner to be $r=1000$. This choice of parameters is typical for full-data KRR problems [5, 25], and the parameters are applied systematically over the 20 regression and classification problems without additional tuning. Note the smaller value of $\mu$ makes many of these problems highly ill-conditioned. We run 250 CG iterations and declare each KRR instance to be “solved” as soon as the relative residual $\lVert(\bm{A}+\mu\mathbf{I})\bm{\beta}-\bm{y}\rVert/\lVert\bm{y}\rVert$ falls below an error tolerance of $\varepsilon=10^{-3}$. This choice of $\varepsilon$ is justified by the fact that the test error plateaus after reaching this tolerance for all the problems we considered.
Larger regularization: $\mu=10^{-7}N$
Smaller regularization: $\mu=10^{-8}N$
@@
Figure 1: Fraction of solved problems versus number of CG iterations for the 20 KRR problem instances in Table˜1.
@@
Figure˜1 charts the fraction of solved problems using RPCholesky and four other preconditioners based on low-rank approximation. Two alternative preconditioners use the partial Cholesky decomposition with greedy pivot selection [26, 57] or uniformly random pivot selection [17, 25]. The final two preconditioners generate randomized low-rank approximations with ridge leverage score sampling (RLS, [1, 39, 48]) or random Fourier features (RFF, [17, 5]). For details of these methods, see Section˜3.1.
@@
RPCholesky is the top-performing method in Figure˜1 with both the larger (left) and the smaller (right) regularization parameters. The closest competitor is RLS, but RPCholesky consistently performs as well or better than RLS. When $\mu/N=10^{-7}$, RPCholesky converges more quickly than RLS on 19 of the 20 test problems and it achieves a speed-up factor of $1.5\times$ for three problems. As another advantage, RPCholesky is simpler to implement and faster to run than RLS, as it eliminates the cost of approximating the ridge leverage scores (see Section˜3.1).
@@
We might worry that RPCholesky changes behavior depending on the random seed, but the algorithm empirically produces results with just $\pm 10\%$ random variation in the number of CG iterations to convergence. To illustrate this point, Figure˜2 shows the distribution of the relative residual $\lVert(\bm{A}+\mu\mathbf{I})\bm{\beta}-\bm{y}\rVert/\lVert\bm{y}\rVert$ for the two problems with the fastest (COMET_MC_SAMPLE) and the slowest (w8a) CG convergence. Lines indicate the median error and shaded regions indicate the $20\%$ to $80\%$ error quantiles, which are computed over 100 independent runs. The error quantiles are hardly distinguishable from the median lines, and they are even closer to the median for RPCholesky than the other stochastic methods (uniform, RLS, and RFF).
@@
w8a Refer to caption
Figure 2: Relative residual versus number of CG iterations for the problem with the fastest (COMET_MC_SAMPLE, left) and slowest (w8a, right) CG convergence when $\mu/N=10^{-7}$. Note the different vertical axis scales.
The importance of eigenvalue decay is also evident in Figure˜2. In the problem with rapid eigenvalue decay (COMET_MC_SAMPLE, left), RPCholesky converges to the accuracy threshold $\varepsilon=10^{-3}$ in just a single iteration. In the problem with slow eigenvalue decay (w8a, right), RPCholesky converges $120\times$ more slowly. All the methods based on low-rank approximation struggle to control the condition number for the w8a problem. Nonetheless, even for this difficult problem, RPCholesky significantly improves on unpreconditioned CG.